OpenStudy (anonymous):

Solve: 1/(x^2-7x+10)=x/(x-5)+1/(x-2)

OpenStudy (anonymous):

This one looks a little scary =S @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Don't worry, it's more afraid of you than you are of it lol

OpenStudy (anonymous):

Haha, I hope so =)

jimthompson5910 (jim_thompson5910):

1/(x^2-7x+10)=x/(x-5)+1/(x-2) 1/((x-5)(x-2))=x/(x-5)+1/(x-2) (x-5)(x-2)[1/((x-5)(x-2))] =(x-5)(x-2)[x/(x-5)] + (x-5)(x-2)[1/(x-2)] 1 = x(x-2) + x-5 Do you see how I got this?

OpenStudy (anonymous):

I think so =) Did you foil x^2-7x+10, then multiply that number by both sides?

jimthompson5910 (jim_thompson5910):

I factored x^2-7x+10 to get (x-5)(x-2), then I multiplied everything by the LCD (x-5)(x-2) to clear out the fractions

OpenStudy (anonymous):

I see! =)

jimthompson5910 (jim_thompson5910):

So now, we do what next?

OpenStudy (anonymous):

simplify to x^2-x-6=0

jimthompson5910 (jim_thompson5910):

good, you got it, then what?

OpenStudy (anonymous):

Plug into quadratic formula

jimthompson5910 (jim_thompson5910):

bingo

OpenStudy (anonymous):

I got -2 and 3

jimthompson5910 (jim_thompson5910):

you nailed it once again

OpenStudy (anonymous):

Yes! Thank you so much!! =)

jimthompson5910 (jim_thompson5910):

as a general rule of thumb, it's best to always check your answers, but they both work (as I just checked them)